# Impact of evanescent scattering modes and finite dispersion in superconducting junctions

### v1

@@ Abstract
Superconducting junctions are essential building blocks for quantum hardware, and their fundamental behavior remains a highly active research field. The behaviour of generic junctions is conveniently described by Beenakker’s determinant formula, linking the subgap energy spectrum to the scattering matrix characterizing the junction. In particular, the gap closing between bound and continuum states in short junctions follows from unitarity of the scattering matrix, and thus, from probability conservation. In this work, we critically reassess two assumptions: that scattering in short junctions is approximately energy-independent and dominated by planar channels. We argue that strongly energy-dependent scattering follows from finite dispersion of the conductor electrons even when they spend little time within the scattering region, and show that evanescent modes play a central role when cross-channel scattering is important. By generalizing Beenakker’s equation and performing a mapping to an effective Hamiltonian, we show that the gap closing is linked to a chiral symmetry. While finite energy-dependence in the scattering breaks the chiral symmetry, we show two distinct mechanisms preserving the gap closing, each connected to new types of constraints on energy-dependent scattering matrices beyond unitarity. If the dispersive mode is planar, the gap closing is still preserved through a time-dependent probability conservation analysis of the scattering process. If the dispersive channel is evanescent, we derive a constraint which, notably, cannot follow from probability conservation. We thus demonstrate that Andreev physics reveal a much wider variety of properties of normal metal scattering than commonly expected. We expect that our findings will have an impact on the dissipative behavior of driven junctions, and offer a new perspective on fundamental properties of scattering matrices.
@@ I Introduction
Superconducting junctions are both of fundamental interest and a key circuit element for quantum hardware. In particular, weak links where one or many channels have a high transparency have been a focal point of research activity in the past decades both theoretically [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32] and experimentally [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. The ground state of a superconductor consists of a coherent many body condensate, out of which single quasi-particles can only be excited above the gap energy $\Delta$, where states exist as an incoherent, delocalized continuum. Yet, if superconductors are linked by a non-superconducting bridge, coherent single-particle excitations can occupy energy levels below the gap, in the form of bound quasi-particle states, so-called Andreev bound states (ABS). As Beenakker showed [1], the formation of these bound states can be directly linked to the scattering properties of the junction. This formalism allows for a convenient and straightforward description of a large variety of systems, such as atomic break junctions [46, 47, 48, 49, 50, 51, 34, 52, 38, 53], multiterminal devices [12, 13, 17, 21, 23, 24, 25, 26, 22, 27, 43, 44], and heterostructures involving topological materials and superconductors [11, 54, 27, 43, 55, 44, 45, 56], to name a few. On a more formal level, the phenomenology of Andreev bound states can be connected to fundamental scattering properties, such as probability conservation (guaranteeing a gap closing between sub- and supergap states in short junctions) or random matrix theory [57, 58, 59, 60], impacting, e.g., the transport properties of chaotic cavities [61, 62, 63], the AC response of short diffusive junctions [17], or the statistical probability to find Weyl points in multiterminal junctions [21].
@@
When the dominant dispersion comes from a regular planar channel, the gap closing is restored due to finite normal backscattering at the superconductor-normal metal (SN) interface, in conjunction with an energy-dependent constraint on the scattering processes which is derived through time-resolved probability conservation considerations. If instead the dispersive channel is evanescent, the gap is closed through regular Andreev reflection – but between evanescent modes. Here, we show that the resulting scattering matrix constraint can no longer be derived from probability conservation, due to the generic impossibility to normalize wave function ansatzes containing diverging evanescent modes. Overall, we thus show that Andreev bound state physics are capable to unravel nontrivial and unexpected constraints on energy-dependent scattering matrices. We explicitly demonstrate the validity of our general concepts and constraints using the example of a ballistic L-shaped junction, a geometry of high experimental relevance [43, 44, 45].
@@
Almost all existing analytic calculations describe the subgap transport at an ideal SN-interface within the so-called Andreev approximation [68, 69], where finite dispersion within the conductor electrons is neglected, resulting in pure Andreev reflection (ideal conversion from electron to hole and vice versa). While there is precedence for going beyond the Andreev approximation [29] (where a finite dispersion leads to a nonzero normal reflection process for planar modes), we find that it is crucial to additionally include resulting strong energy-dependencies in the normal scattering matrix $S$, since the resulting ABS spectrum near the gap is highly sensitive to the interplay of both SN-reflection and scattering. Moreover, the behavior of evanescent modes at the SN-interface has not been considered at all in the existing literature. In this work, we show that finite dispersion renders the evanescent analog of Andreev reflection nonzero, which is at the origin of the above summarized gap closing phenomenon, and resulting nontrivial scattering matrix contraints.
@@
Overall, our work generalizes the state of the art $S$-matrix formalism (Beenakker framework) as a powerful and convenient tool to analyse ABS physics for generic multiterminal and multichannel junctions. We expect that the here uncovered spectral properties close to the superconducting gap are of importance for future projects to improve our understanding of dissipative properties of driven weak links (building on, e.g., Refs. [8, 10]). Finally, the explicit inclusion of evanescent modes in the scattering problem (and our demonstration of their importance) might provide new impulses within the field of random matrix theory [57, 58, 59]: namely, as of now, the literature only describes the distribution of scattering coefficients for planar waves (e.g., the Dorotkov distribution for short diffusive wires [60]), whereas the distribution of evanescent coefficients has to the best of our knowledge not been studied. Moreover, the understanding of constraints relating scattering matrices at different energies could potentially even provide an alternative, simpler access to the prediction of probability distributions of scattering coefficients as a function of energy.
@@
This work is organized as follows. Section II recapitulates the standard Beenakker formalism. In Section III the impact of finite dispersion on energy dependent scattering and ABS is explored. Section IV introduces the revised Beenakker framework with evanescent modes and energy dependent scattering fully included. In Section V evanescent scattering and Andreev reflection is related to a constraint on the scattering matrix. In Section VI the revised framework is applied to a junction with non-trivial geometry and the effects on the ABS spectra are verified by explicit numerical calculation.
@@ II Summary of standard Beenakker formalism and main results
This section offers an overview of the formation of Andreev bound states in superconducting junctions and it’s currently accepted treatment including commonly applied approximations. At the end of the section, we summarize our main findings and accomplishments.
Let us consider a generic junction with $M$ superconducting terminals with a superconducting gap $\Delta$. Let them be joined by a central region consisting either of a normal metal or semiconductor material. Electrons and holes are assumed to be subject to elastic scattering within the central region. Focusing on energies below $\Delta$, electrons and holes undergo Andreev reflection [68] at the interface to the superconductor (SN-interface).
@@
where the vector $\psi^{\text{in}}$ ($\psi^{\mathrm{out}}$) denotes the amplitudes of modes incident (outgoing) with respect to the scattering center. The scattering matrix $S$ is block diagonal
@@
The finite cross section area of the conductor arms joining at the center of the normal region leads to the formation of channels. Assuming ballistic transport in these arms, their physics can be described by the dispersion relation
@@
where $T_{\mathrm{\mu,\nu}}^{m,n}$ denotes transmission from the $\nu$th to the $\mu$th terminal. Note that we here for simplicity assume that each terminal has the same number of channels (symmetric conductor arms). The formalism is however very easily generalized to the case where different terminals may have different channel numbers.
At this stage it is important to point out, that in the present work, we choose to stick to a slightly different definition of the multi-channel scattering matrix. In the ansatz of the wave function for the multi-channel regime, it is common practice [70] to add a normalization prefactor related to the group velocity, ${\sim\sqrt{v_{n}}}$ in front of the wave function component of channel $n$. For our purposes, we find it more convenient to omit this prefactor, as it simplifies some analytic proofs and the numerical computations below. Moreover, such a normalization choice seems even more impractical with the inclusion of evanescent modes (virtual occupations of higher energy channels with imaginary $k$), as they have no meaningful definition of a group velocity. At any rate, for the purely planar mode description, our scattering matrix $S$ is related to the more commonly used literature definition $\widetilde{S}$ through ${\widetilde{S}=\sqrt{\mathbf{v}}S\sqrt{\mathbf{v}}^{-1}}$, where $\mathbf{v}$ is a diagonal matrix with the group velocities of the corresponding channels as its entries. Note that for the single-channel case, $\widetilde{S}=S$, such that this distinction is only important for multi-channel scattering problems (with distinct group velocities). While $\widetilde{S}$ is unitary due to probability conservation, the matrix $S$ satisfies $S^{\dagger}\mathbf{v}S=\mathbf{v}$. Similarly, the presence of time-reversal symmetry (TRS), $\widetilde{S}^{T}=\widetilde{S}$ here manifests as $S^{T}\mathbf{v}=\mathbf{v}S$. Some important identities however remain the same, such as the combined properties of probability conservation and TRS, yielding $\widetilde{S}^{*}\widetilde{S}=S^{*}S=1$.
@@
and scattering reduces to a set of independent ${1\mathrm{D}}$ problems. As a consequence, there is no longer any cross-channel scattering, ${S_{e,h}^{m,n}=0}$ for ${n\neq m}$ (also meaning that here, ${\widetilde{S}=S}$). Conversely, dropping this separability assumption leads to interactions between different channels. Cross-channel scattering will play an integral part in this work, as we detail further below.
@@
where $\phi_{l}$ is the phase of the $l$th superconducting terminal. Under the above assumptions, the reflection coefficient $\alpha$ is identical for each channel ${n\leq n_{\mathrm{P}}}$ involved in the process, such that we only need the unit matrix $\mathds{1}_{n_{\mathrm{P}}\times n_{\mathrm{P}}}$ (here, actually spanning across the number of planar channels $n_{\mathrm{P}}$, which will become important later). Assuming that the spatial dependence of the superconducting pairing is mainly along the propagation direction ($x$), Andreev reflection can be treated for each channel independently, in the same spirit as for the discussion after Eq. (8). For simplicity, we shall continue to uphold this assumption throughout this work. Consequently, unlike for normal scattering, cross-channel interactions will not be incorporated in $R$. As a consequence, our previously discussed choice to not include group velocities explicitly in the ansatz of the wave functions would have no impact here, and our definition of $R$ remains consistent with the standard literature definition. Note that matters are different, once we go beyond the Andreev approximation. Here, electrons and holes have (slightly) different group velocities, such that here, different normalization conventions could impact the final form of $R$. Again, we avoid such complications by not performing a normalization with respect to group velocities altogether.
Figure 1: Importance of evanescent modes for the bound states spectrum in short vs. long conductor arms. The figure depicts a conductor arm (light gray region) connecting the scattering center (dark gray region) to one of the superconducting terminals (blue region). In the most general case, the solution of the central scattering problem involves (virtually occupied) evanescent modes due to energetically forbidden channels. Panel a): For long conductor arms, conventional Andreev reflection at the SN-interface and normal scattering at the center result in planar standing wave configurations of electron and hole excitations (green). Potential evanescent modes exiting the scattering region (dark blue real exponential) decay before reaching the SN-interface. Panel b): For short conductor arms, evanescent modes penetrate the SN-interface such that evanescent Andreev reflection becomes important. In this situation, the standing waves in the conductor arms form superpositions consisting of both decaying and diverging real exponentials, in addition to planar modes.
From the perspective of single particle excitations within the normal lead, the physics of tunnel junctions below the superconducting gap closely resembles that of a particle confined in a box potential. The boundary conditions, as defined in Eqs. (1) and (10), describe a scenario where electrons undergo Andreev reflection into holes at the SN-interfaces, which are then scattered at the center, and vice versa. In order to maintain a standing wave configuration, this Andreev loop [71, 72] needs to create a coherent superposition of particle and hole excitations. By combining the boundary conditions, Eqs. (1) and (10), the corresponding interference condition can be formulated as the eigenvalue problem [1]
@@
where both scattering and Andreev-reflection is energy dependent ($S\left[E\right],\ R\left[E\right]$) in general. If $R$ is of the form given in Eq. (11), the interference condition assumes the form
@@
To simplify this equation even further, one usually invokes the notion of a short junction limit. In the existing literature, this limit is characterized by the Thouless energy, which is in turn defined by the inverse dwelling time ${E_{\mathrm{Th}}\equiv 1/\tau_{\mathrm{dw}}}$. The dwelling time ${\tau_{\mathrm{dw}}}$ represents the time scale the particle resides within the scattering center. For a ballistic scattering region, this time can be estimated as ${\tau_{\mathrm{dw}}\sim v_{F}/L}$ where $L$ indicates the size of the region and $v_{F}$ is the Fermi velocity.
@@
In this work, we explicitly extend the framework to account for evanescent modes and identify specific conditions under which they become of importance. We find in particular that when one or several of these channels have finite dispersion, they cause two related effects which both impact the ABS spectrum. First, finite dispersion in conjunction with cross-channel coupling yields a strong energy dependence of the scattering matrix, notably even if the Thouless energy is large. Secondly, Andreev reflection of evanescent channels becomes particularly relevant, such that not only $S$, but also $R$, needs to be extended to accommodate evanescent modes. While strong energy-dependence of $S$ alone would provide a detaching of the bound states from the continuum (as it invalidates the conditions leading to Eq. (20) above), the evanescent Andreev reflection allows to exactly compensate this detachment. Curiously, this special combination of mechanisms culminates in a formulation of energy-dependent constraints on the full scattering matrix (including planar and evanescent modes) which, crucially, cannot be derived from probability conservation arguments. To summarize, our main achievements are:
@@
The above points are developed in detail in the following sections. After developing some general findings regarding the ABS spectrum and the energy dependence of the normal metal scattering process, we setup the new formalism, and subsequently apply it to a specific, experimentally relevant junction geometry.
III Impact of energy-dependent scattering on ABS spectrum

### v2

# Interplay between evanescent scattering modes and finite dispersion in superconducting junctions

@@ Abstract
Superconducting junctions are essential building blocks for quantum hardware, and their fundamental behaviour remains a highly active research field. The behaviour of generic junctions is conveniently described by Beenakker’s determinant formula, linking the subgap energy spectrum to the scattering matrix characterising the junction. In particular, the gap closing between bound and continuum states in short junctions follows from unitarity of the scattering matrix, and thus, from probability conservation. In this work, we critically reassess two assumptions: that scattering in short junctions is approximately energy-independent and dominated by planar channels. We argue that strongly energy-dependant scattering follows from finite dispersion of the conductor electrons even when they spend little time within the scattering region, and show that evanescent modes play a central role when cross-channel scattering is important. By generalising Beenakker’s equation and performing a mapping to an effective Hamiltonian, we show that the gap closing is linked to a chiral symmetry. While finite energy dependence in the scattering breaks the chiral symmetry, we show two distinct mechanisms preserving the gap closing, each connected to new types of constraints on energy-dependant scattering matrices beyond unitarity. If the dispersive mode is planar, the gap closing is still preserved through a time-dependant probability conservation analysis of the scattering process. If the dispersive channel is evanescent, we derive a constraint which, notably, cannot follow from probability conservation. We thus demonstrate that Andreev physics reveal a much wider variety of properties of normal-metal scattering than commonly expected. We expect that our findings will have an impact on the dissipative behaviour of driven junctions, and offer a new perspective on fundamental properties of scattering matrices.
@@ I Introduction
Superconducting junctions are both of fundamental interest and a key circuit element for quantum hardware. In particular, weak links where one or many channels have a high transparency have been a focal point of research activity in the past decades both theoretically [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] and experimentally [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. The ground state of a superconductor consists of a coherent many-body condensate, out of which single quasiparticles can only be excited above the gap energy $\Delta$, where states exist as an incoherent, delocalised continuum. Yet, if superconductors are linked by a non-superconducting bridge, coherent single-particle excitations can occupy energy levels below the gap, in the form of bound quasiparticle states, so-called Andreev bound states (ABS). As Beenakker showed [1], the formation of these bound states can be directly linked to the scattering properties of the junction. This formalism allows for a convenient and straightforward description of a large variety of systems, such as atomic break junctions [47, 48, 49, 50, 51, 52, 35, 53, 39, 54], multiterminal devices [12, 13, 17, 21, 23, 24, 25, 26, 22, 27, 44, 45], and heterostructures involving topological materials and superconductors [11, 55, 27, 44, 56, 45, 46, 57], to name a few. On a more formal level, the phenomenology of Andreev bound states can be connected to fundamental scattering properties, such as probability conservation (guaranteeing a gap closing between sub- and supergap states in short junctions) or random matrix theory [58, 59, 60, 61], impacting, e.g., the transport properties of chaotic cavities [62, 63, 64], the AC response of short diffusive junctions [17], or the statistical probability to find Weyl points in multiterminal junctions [21].
@@
When the dominant dispersion comes from a regular planar channel, the gap closing is restored due to finite normal backscattering at the superconductor-normal metal (SN) interface, in conjunction with an energy-dependant constraint on the scattering processes which is derived through time-resolved probability conservation considerations. If instead the dispersive channel is evanescent, the gap is closed through regular Andreev reflection – but between evanescent modes. Here, we show that the resulting scattering matrix constraint can no longer be derived from probability conservation, due to the generic impossibility to normalise wave function ansatzes containing diverging evanescent modes. Overall, we thus show that Andreev bound state physics are capable to unravel nontrivial and unexpected constraints on energy-dependant scattering matrices. We explicitly demonstrate the validity of our general concepts and constraints using the example of a ballistic L-shaped junction, a geometry of high experimental relevance [44, 45, 46].
@@
Almost all existing analytic calculations describe the subgap transport at an ideal SN-interface within the so-called Andreev approximation [68, 69], where finite dispersion within the conductor electrons is neglected, resulting in pure Andreev reflection (ideal conversion from electron to hole and vice versa). While there is precedence for going beyond the Andreev approximation [29] (where a finite dispersion leads to a nonzero normal reflection process for planar modes), we find that it is crucial to additionally include resulting strong energy-dependencies in the normal scattering matrix $S$, since the resulting ABS spectrum near the gap is highly sensitive to the interplay of both SN-reflection and scattering. Moreover, the behaviour of evanescent modes at the SN-interface has not been considered at all in the existing literature. In this work, we show that finite dispersion renders the evanescent analogue of Andreev reflection nonzero, which is at the origin of the above summarised gap closing phenomenon, and resulting nontrivial scattering matrix constraints.
@@
Overall, our work generalises the state of the art $S$-matrix formalism (Beenakker framework) as a powerful and convenient tool to analyse ABS physics for generic multiterminal and multichannel junctions. We expect that the here uncovered spectral properties close to the superconducting gap are of importance for future projects to improve our understanding of dissipative properties of driven weak links (building on, e.g., Refs. [8, 10]). Finally, the explicit inclusion of evanescent modes in the scattering problem (and our demonstration of their importance) might provide new impulses within the field of random matrix theory [58, 59, 60]: namely, as of now, the literature only describes the distribution of scattering coefficients for planar waves (e.g., the Dorotkov distribution for short diffusive wires [61]), whereas the distribution of evanescent coefficients has to the best of our knowledge not been studied. Moreover, the understanding of constraints relating scattering matrices at different energies could potentially even provide an alternative, simpler access to the prediction of probability distributions of scattering coefficients as a function of energy.
@@
This work is organised as follows. Section II recapitulates the standard Beenakker formalism. In Section III the impact of finite dispersion on energy dependant scattering and ABS is explored. Section IV introduces the revised Beenakker framework with evanescent modes and energy dependant scattering fully included. In Section V evanescent scattering and Andreev reflection is related to a constraint on the scattering matrix. In Section VI the revised framework is applied to a junction with non-trivial geometry and the effects on the ABS spectra are verified by explicit numerical calculation.
@@ II Summary of standard Beenakker formalism and main results
Figure 1: Importance of evanescent modes for the bound states spectrum in short vs. long conductor arms. The figure depicts a conductor arm (light grey region) connecting the scattering centre (dark grey region) to one of the superconducting terminals (blue region). In the most general case, the solution of the central scattering problem involves (virtually occupied) evanescent modes due to energetically forbidden channels. Panel a): For long conductor arms, conventional Andreev reflection at the SN-interface and normal scattering at the centre result in planar standing wave configurations of electron and hole excitations (green). Potential evanescent modes exiting the scattering region (dark blue real exponential) decay before reaching the SN-interface. Panel b): For short conductor arms, evanescent modes penetrate the SN-interface such that evanescent Andreev reflection becomes important. In this situation, the standing waves in the conductor arms form superpositions consisting of both decaying and diverging real exponentials, in addition to planar modes.
To provide an overview of the standard $S$-matrix formalism, let us consider a generic junction with $M$ superconducting terminals with a superconducting gap $\Delta$. Let them be joined by a central region consisting either of a normal metal or a semiconductor material (Fig. 1 depicts a conductor arm connecting the scattering centre with one of the superconducting terminals). Electrons and holes are assumed to be subject to elastic scattering within the central region. Focusing on energies below $\Delta$, electrons and holes undergo Andreev reflection [68] at the interface to the superconductor (SN-interface).
@@
where the vector $\psi^{\text{in}}$ ($\psi^{\mathrm{out}}$) denotes the amplitudes of modes incident (outgoing) with respect to the scattering centre. The scattering matrix $S$ is block diagonal
@@
The finite cross section area of the conductor arms joining at the centre of the normal region leads to the formation of channels. Assuming ballistic transport in these arms, their physics can be described by the dispersion relation
@@
where $T_{\mathrm{\mu,\nu}}^{m,n}$ denotes transmission from the $\nu$th to the $\mu$th terminal. Note that we assume here for simplicity that each terminal has the same number of channels (symmetric conductor arms). The formalism is however very easily generalised to the case where different terminals may have different channel numbers.
At this stage it is important to point out, that in the present work, we choose to stick to a slightly different definition of the multi-channel scattering matrix. In the ansatz of the wave function for the multi-channel regime, it is common practice [70] to add a normalisation prefactor related to the group velocity, ${\sim\sqrt{v_{n}}}$ in front of the wave function component of channel $n$. For our purposes, we find it more convenient to omit this prefactor, as it simplifies some analytic proofs and the numerical computations below. Moreover, such a normalisation choice seems even more impractical with the inclusion of evanescent modes (virtual occupations of higher energy channels with imaginary $k$), as they have no meaningful definition of a group velocity. At any rate, for the purely planar mode description, our scattering matrix $S$ is related to the more commonly used literature definition $\widetilde{S}$ through ${\widetilde{S}=\sqrt{\mathbf{v}}S\sqrt{\mathbf{v}}^{-1}}$, where $\mathbf{v}$ is a diagonal matrix with the group velocities of the corresponding channels as its entries. Note that for the single-channel case, $\widetilde{S}=S$, so this distinction is only important for multi-channel scattering problems (with distinct group velocities). While $\widetilde{S}$ is unitary due to probability conservation, the matrix $S$ satisfies $S^{\dagger}\mathbf{v}S=\mathbf{v}$. Similarly, the presence of time-reversal symmetry (TRS), $\widetilde{S}^{T}=\widetilde{S}$ here manifests as $S^{T}\mathbf{v}=\mathbf{v}S$. Some important identities however remain the same, such as the combined properties of probability conservation and TRS, yielding $\widetilde{S}^{*}\widetilde{S}=S^{*}S=1$.
@@
and scattering reduces to a set of independent one-dimensional (${1\mathrm{D}}$) problems. As a consequence, there is no longer any cross-channel scattering, ${S_{e,h}^{m,n}=0}$ for ${n\neq m}$ (also meaning that here, ${\widetilde{S}=S}$). Conversely, dropping this separability assumption leads to interactions between different channels. Cross-channel scattering will play an integral part in this work, as we detail further below.
@@
where $\phi_{l}$ is the phase of the $l$th superconducting terminal. Under the above assumptions, the reflection coefficient $\alpha$ is identical for each channel ${n\leq n_{\mathrm{P}}}$ involved in the process, such that we only need the unit matrix $\mathds{1}_{n_{\mathrm{P}}\times n_{\mathrm{P}}}$ (here, actually spanning across the number of planar channels $n_{\mathrm{P}}$, which will become important later). Assuming that the spatial dependence of the superconducting pairing is mainly along the propagation direction ($x$), Andreev reflection can be treated for each channel independently, in the same spirit as for the discussion after Eq. (8). For simplicity, we continue to uphold this assumption throughout this work. Consequently, unlike for normal scattering, cross-channel interactions will not be incorporated in $R$. As a consequence, our previously discussed choice to not include group velocities explicitly in the ansatz of the wave functions would have no impact here, and our definition of $R$ remains consistent with the standard literature definition. Note that matters are different, once we go beyond the Andreev approximation. Here, electrons and holes have (slightly) different group velocities, such that here, different normalisation conventions could impact the final form of $R$. Again, we avoid such complications by not performing a normalisation with respect to group velocities altogether.
From the perspective of single particle excitations within the normal lead, the physics of tunnel junctions below the superconducting gap closely resembles that of a particle confined in a box potential. The boundary conditions, as defined in Eqs. (1) and (10), describe a scenario where electrons undergo Andreev reflection into holes at the SN-interfaces, which are then scattered at the centre, and vice versa. In order to maintain a standing wave configuration, this Andreev loop [71, 72] needs to create a coherent superposition of particle and hole excitations. By combining the boundary conditions, Eqs. (1) and (10), the corresponding interference condition can be formulated as the eigenvalue problem [1]
@@
where both scattering and Andreev-reflection is energy dependant ($S\left[E\right],\ R\left[E\right]$) in general. If $R$ is of the form given in Eq. (11), the interference condition assumes the form
@@
To simplify this equation even further, one usually invokes the notion of a short junction limit. In the existing literature, this limit is characterised by the Thouless energy, which is in turn defined by the inverse dwelling time ${E_{\mathrm{Th}}\equiv 1/\tau_{\mathrm{dw}}}$. The dwelling time ${\tau_{\mathrm{dw}}}$ represents the time scale the particle resides within the scattering centre. For a ballistic scattering region1, this time can be estimated as ${\tau_{\mathrm{dw}}\sim v_{F}/L}$ where $L$ indicates the size of the region and $v_{F}$ is the Fermi velocity.
@@
In this work, we explicitly extend the framework to account for evanescent modes and identify specific conditions under which they become of importance. We find in particular that when one or several of these channels have finite dispersion, they cause two related effects which both impact the ABS spectrum. First, finite dispersion in conjunction with cross-channel coupling yields a strong energy dependence of the scattering matrix, notably even if the Thouless energy is large. Secondly, Andreev reflection of evanescent channels becomes particularly relevant, such that not only $S$, but also $R$, needs to be extended to accommodate evanescent modes. While strong energy-dependence of $S$ alone would provide a detaching of the bound states from the continuum (as it invalidates the conditions leading to Eq. (20) above), the evanescent Andreev reflection allows to exactly compensate this detachment. Curiously, this special combination of mechanisms culminates in a formulation of energy-dependant constraints on the full scattering matrix (including planar and evanescent modes) which, crucially, cannot be derived from probability conservation arguments. To summarise, our main achievements are that
@@
The above points are developed in detail in the following sections. After developing some general findings regarding the ABS spectrum and the energy dependence of the normal metal scattering process (cf. Sec. III), we setup the new formalism (cf. Sec. IV), and subsequently apply it to a specific, experimentally relevant junction geometry (Sec. VI).
III Impact of energy-dependant scattering on Andreev bound-state spectrum